Math 154a: Statistics

Robert Gould  Office: MS 6151 x6-3381  Office Hours:M 10-11, W 1-2

Purpose of Course

Statistics is a discipline employed by a wide variety of people in a variety of professions. Briefly, it is the study of the use of uncertainty to analyze data. Uncertainty occurs almost everywhere, from measurement errors, to unpredictable cardgames, to unknowable reactions to drugs. Part of the job of the statistician is to see past this noise of randomness to detect and observe the underlying phenomenon. Thus, statistics is used to answer questions such as Do power lines cause cancer? Does the DNA sample match the suspect’s? Are SAT scores a good predictor of collegiate success?

Statistics rests solidly on a mathematical foundation composed mostly of probability, and so we’ll be spending a substantial part of this course on probability theory. This is a mathematics course, and the prerequisites demand a background in multivariable calculus, so we will be treating ”mathematical issues.” But we will also keep an eye on applications, and be concerned about good and bad applications of statistics (although honestly the bad applications are far more entertaining.) A word of warning: it will be at least five weeks before we can begin to apply our mathematical foundation to some actual data.

The ability to clearly communicate your ideas and solutions is absolutely necessary to your success in mathematics and the industrial world. Forcing yourself to think about how you will organize your solution can also assist you in solving a problem. For this reason, we will emphasize clear, written solutions, as well as correct answers.

Prerequisites The prerequisites for this course are Math 32B (Calculus of Several Variables) and Math 33B (Infinite Series).

Computer You have been given a computer account. We will be making extensive and dare I say novel use of the computer. Please see the handout on the World Wide Web.

Materials You need the book An Introduction to Mathematical Statistics and its Applications, by Richard J. Larsen and Morris L. Marx. Also, all students are required to buy a bluebook to use as a “journal”, in which you’ll record certain computer exercises. If you own a computer and a modem, you will find BruinOnline extremely helpful, but this is not required. BruinOnline is sold at the ASUCLA store for about $29. You will also need a calculator, but nothing fancy is required.

Class Discussion I believe that discussion is an integral part of the learning process. It assists me by letting me know where you are having difficulties, and its helpful for you to work things out outloud. There are three opportunities for discussion: in lecture, (asking or answering questions), in discussion section (asking questions or helping fellow students with questions), or on the computer (more about this later.) There is no easy way to grade discussion, (except for the computer discussion), so my policy is to use class participation as a way of deciding borderline grades.

Grades There will be two midterms and a final. In addition, you have weekly homework assignments and two “essays”. You will also be graded on your use of the computer “Chat Room”, a feature that is described on the other handout.

1. Chatroom If you don’t participate, you will be penalized.

2. Homework, Computer essays 10% + 10%

3. Midterm 1 Friday, October 27. 20%

4. Midterm 2 Friday, November 17. 20%

5. Final Wednesday, December 13, 8 am. 40%

Late/makeup Policy There will be no make-up exams. Late homework will not be accepted. Please write these dates down and turn in your homework on time so that there will be no problems.

Homework Homework will be assigned every Friday and due in class the following Friday. Your homework should be clearly labelled with your name, student ID number, section number, and the date. Problems should be clearly identified. You are responsible for providing a clear and lucid explanation of your work, as well as an answer! If the grader has trouble understanding you, or even worse, cannot read your writing, then he or she has the option of giving you a 0. You are encouraged to work with others, but the work you turn in must be your own.

Computer Essays Each week, an essay topic will be assigned via the 154 Homepage on the computer. These topics will typically be problems with interesting backgrounds, or famous unintuitive problems, or sometimes problems requiring some data analysis. You should record these problems in your bluebook and your thoughts about them. Use the Chat Room to talk about your ideas and see what your classmates think. At the end of the quarter, you must submit essays on two of the topics. Details on these essays will be provided later, but each essay should be 1 to 2 pages long, and should provide a discussion of the problem and a description of how your reached an answer. You might want to select discussions appearing in the Chat Room to support your argument and/or to criticize.

Midterms and Final The exams in this class are all loosely based on, perhaps “inspired by” is a better phrase, the homework. If you understand the homework, then you are prepared for the exams. Do not expect homework problems to appear verbatim, however. Typically, there will be one or two more challenging problems on each exam designed to see how you can apply the concepts to a novel situation.